OpenStudy (anonymous):

Kala is going to rent a truck for one day. There are two companies she can choose from, and they have the following prices. Company A charges $102 and allows unlimited mileage Company B has an initial fee of $75 and charges an additional $0.60 for every mile driven. m For what mileages will Company A charge less than Company B? Use m for the number of miles driven, and solve your inequality for m.

OpenStudy (anonymous):

\[102 \le 75 + 0.6(miles)\]\[\frac{102-75}{0.6}\le x (miles)\]\[45\le (miles)\]

OpenStudy (anonymous):

thnks :)

OpenStudy (texaschic101):

102 = 75 + .60m 102 - 75 = .60m 27 = .60m 27/.60 = m 45 = m at 45 miles they will charge equal amounts. So for company A to charge less, they would have to drive more then 45 miles. (46 and above)

OpenStudy (texaschic101):

let me check something......hold on

OpenStudy (texaschic101):

102 < 75 + .6(46) 102 < 75 + 27.60 102 < 102.60 102 < 75 + .6(47) 102 < 75 + 28.20 102 < 103.20 thats what I thought...for company A to be cheaper, the person would have to drive 46 or MORE miles

OpenStudy (anonymous):

thanks so much :)

OpenStudy (texaschic101):

no problem :)